the domain of function:


f(x) = sin-1[(2x-1)/(3x-2)] + cos-1[(3x-4)/5]

Grade:11

1 Answers

Shadow
10 years ago
[(2x-1)/(3x-2)>=-1&&(3x-4)/5>-1]&&[(2x-1)/(3x-2)<=1 &&((3x-4)/5)<=1]